Consider three different processors P1, P2, and P3 executing the same instruction set. P1 has a 3 GHz clock rate and a CPI of 1.5. P2 has a 3 GHz clock rate and a CPI of 1.0. P3 has a 2.5 GHz clock rate and has a CPI of 2.2.

a. Which processor has the highest performance expressed in instructions per second?

b. If each processor executes a program in 35 seconds, find the number of cycles and the number of instructions.

c. We are trying to reduce the execution time by 15% but this leads to an increase of 10% in the CPI. What clock rate should we have to get this time reduction?
Answer #1

`Hey,

Note: Brother in case of any queries, just comment in box I would be very happy to assist all your queries

a.

Clock Rate = clock cycles per second.

CPI = Clock cycle per instruction, which is clock cycles required by processor to execute one instruction.

For P1:

Clock Rate = 3GHz = 3*(10^9) clock/second

CPI = 1.5 clock/instruction

Performance (instruction/second) = Clock Rate/CPI

3*(10^9) divide by 1.5 = 2*(10^9)

For P2:

Clock Rate = 3GHz = 3*(10^9) clock/second

CPI = 1 clock/instruction

Performance (instruction/second) = Clock Rate/CPI

3*(10^9) divide by 1 = 3*(10^9)

For P3:

Clock Rate = 2.5GHz =2.5*(10^9) clock/second

CPI = 2.2 clock/instruction

Performance (instruction/second) = Clock Rate/CPI

2.5*(10^9) divide by 2.2 = 1.14*(10^9)

It means P2 is has the highest performance in terms of Instruction per second.

b.

Clock Rate = clock cycles per second.​

For P1 -

Clock Rate = 3GHz = 3*(10^9) "clock cycles"/second

Execution time for program = 35 seconds

Number of Cycles for 35 seconds = 3*(10^9)*35 = 10.5*(10^10)

Number of instructions for 1 seconds = 2*(10^9) (Solved in "a" part)

Number of instructions for 35 seconds = 2*(10^9) * 35 = 7*10^10

For P2 -

Clock Rate = 3GHz = 3*(10^9) "clock cycles"/second

Execution time for program = 35 seconds

Number of Cycles for 35 seconds = 3*(10^9)*35 = 10.5*(10^10)

Number of instructions for 1 seconds = 3*(10^9) (Solved in "a" part)

Number of instructions for 35 seconds = 3*(10^9) * 35 = 10.5*10^10

For P3 -

Clock Rate = 2.5GHz = 2.5*(10^9) "clock cycles"/second

Execution time for program = 35 seconds

Number of Cycles for 35 seconds = 2.5*(10^9)*35 = 8.75*(10^10)

Number of instructions for 1 seconds = 1.14*(10^9) (Solved in "a" part)

Number of instructions for 35 seconds = 1.14*(10^9) * 35 = 3.90*10^10

Note: Brother according to HomeworkLib's policy we are only allowed to answer first 2 part if there are many. So, I request you to post other part as separate posts.

Kindly revert for any queries

Thanks.